# Non-Markovian Reduced Models to Unravel Transitions in Non-equilibrium Systems

### v1

@@ Acknowledgements.
This work has been supported by the Office of Naval Research (ONR) Multidisciplinary University Research Initiative (MURI) grant N00014-20-1-2023, and by the National Science Foundation grants DMS-2108856, DMS-2407483, and DMS-2407484. This work has been also supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 810370) and by a Ben May Center grant for theoretical and/or computational research.
@@
Since the OU process $z_{q}(s,\omega)$ has sublinear growth (see e.g. [29, Lemma 3.1]), that is $\lim_{s\rightarrow\pm\infty}\frac{z_{q}(s,\omega)}{s}=0$ for all $\omega$, one can readily check that the integral $I_{K_{z}}$ defined in (105) is finite for all $t$ provided that
$\Re\left(\lambda_{n}-\sum_{p\in K\setminus K_{z}}\lambda_{j_{p}}\right)<0.$ (106)
@@
$\frac{\partial\mathfrak{J}(t,X,\omega)}{\partial t}$ $=\Pi_{\mathfrak{s}}G_{k}(X+z_{\mathfrak{c}}(\theta_{t}\omega))-A_{\mathfrak{s}}\mathfrak{J}(t,X,\omega)$ (107)
$\quad-\int_{-\infty}^{t}e^{(t-s)A_{\mathfrak{s}}}\Pi_{\mathfrak{s}}DG_{k}\Big{(}e^{(s-t)A_{\mathfrak{c}}}X+z_{\mathfrak{c}}(\theta_{s}\omega)\Big{)}e^{(s-t)A_{\mathfrak{c}}}A_{\mathfrak{c}}X\,\mathrm{d}s,$
@@
$\int_{-\infty}^{t}e^{(t-s)A_{\mathfrak{s}}}\Pi_{\mathfrak{s}}DG_{k}\Big{(}e^{(s-t)A_{\mathfrak{c}}}X+z_{\mathfrak{c}}(\theta_{s}\omega)\Big{)}e^{(s-t)A_{\mathfrak{c}}}A_{\mathfrak{c}}X\,\mathrm{d}s,$
@@
$q(s,\omega)$ $=\int_{t-\tau}^{s}e^{(s-s^{\prime})A_{\mathfrak{s}}}\Pi_{\mathfrak{s}}G_{k}\Big{(}e^{(s^{\prime}-s)A_{\mathfrak{c}}}X+z_{\mathfrak{c}}(s^{\prime},\omega)\Big{)}\,\mathrm{d}s^{\prime}.$
@@
$q_{\tau}(t,X,\omega)=\int_{t-\tau}^{t}e^{(t-s^{\prime})A_{\mathfrak{s}}}\Pi_{\mathfrak{s}}G_{k}\Big{(}e^{(s^{\prime}-t)A_{\mathfrak{c}}}X+z_{\mathfrak{c}}(s^{\prime},\omega)\Big{)}\,\mathrm{d}s^{\prime}.$
@@
$\int_{-\infty}^{t-\tau}e^{(t-s)A_{\mathfrak{s}}}\Pi_{\mathfrak{s}}G_{k}(e^{(s-t)A_{\mathfrak{c}}}X+z_{\mathfrak{c}}(s,\omega))\,\mathrm{d}s.$
The boundedness of $\mathfrak{J}$ implies that the $H_{\mathfrak{s}}$-norm of the integral on the RHS of (109) converges to zero as $\tau$ goes to infinity, and (21) follows.
@@ Appendix B Solving the sACE and its optimal reduced system
For the numerical integration of the stochastic Allen-Cahn equation (74), we adopt a semi-implicit Euler method, where the nonlinearity $-u^{3}$ and the noise term $\,\mathrm{d}{\bm{W}_{t}}$ are treated explicitly, and the linear term $\partial_{x}^{2}u+u$ is treated implicitly; the spatial discretization is handled with a pseudo-spectral method; see [30, Section 6.1] for a detailed description of this scheme applied to a stochastic Burgers-type equation. We have set the time-step size to be $\delta t=10^{-2}$ and the spatial resolution to be $\delta x=L/201$.
@@ C.1 Computation of the double-fold bifurcation diagram
The numerical integration of the stochastic equation (94) is performed in the same way as for the sACE outlined in Appendix B using a semi-implicit Euler method. Here also the nonlinearity and the noise term are treated explicitly, while the linear term is treated implicitly. The spatial discretization is handled with a pseudo-spectral method; see again [30, Section 6.1]. We have set the time-step size to be $\delta t=10^{-2}$ and the spatial resolution to be $\delta x=L/257$. The initial data is set to be $u_{0}=0.5\bm{e}_{1}$ for the results shown in Figs. 12 and 15.
The optimal reduced system (Eq. (LABEL:Eq_reduced_S_shaped)) is integrated with a forward Euler method still with $\delta t=10^{-2}$. The initial condition $J^{n}_{\tau^{*}_{n}}(\bm{\beta},t_{0};\omega)$ (cf. (97)) for integrating (98) is computed using the trapezoidal rule, and $t_{0}=\max\{\tau_{3}^{*},\tau_{5}^{*}\}$. The initial data for Eq. (LABEL:Eq_reduced_S_shaped) is set to be the projection of the SPDE solution onto $\bm{e}_{1}$ at time $t=t_{0}$.

### v2

@@ Acknowledgements.
Several insights discussed in this work have substantially benefited from the reviewers’ constructive comments, and we express our deep gratitude to them. This work has been supported by the Office of Naval Research (ONR) Multidisciplinary University Research Initiative (MURI) grant N00014-20-1-2023, and by the National Science Foundation grants DMS-2108856, DMS-2407483, and DMS-2407484. This work has been also supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 810370) and by a Ben May Center grant for theoretical and/or computational research.
@@
Since the OU process $z_{q}(s,\omega)$ has sublinear growth (see e.g. [30, Lemma 3.1]), that is $\lim_{s\rightarrow\pm\infty}\frac{z_{q}(s,\omega)}{s}=0$ for all $\omega$, one can readily check that the integral $I_{K_{z}}$ defined in (111) is finite for all $t$ provided that
$\mathrm{Re}\left(\lambda_{n}-\sum_{p\in K\setminus K_{z}}\lambda_{j_{p}}\right)<0.$ (112)
@@
$\frac{\partial\mathfrak{J}(t,X,\omega)}{\partial t}$ $=\Pi_{\mathfrak{f}}G_{k}(X+z_{\mathfrak{c}}(\theta_{t}\omega))-A_{\mathfrak{f}}\mathfrak{J}(t,X,\omega)$ (113)
$\quad-\int_{-\infty}^{t}e^{(t-s)A_{\mathfrak{f}}}\Pi_{\mathfrak{f}}DG_{k}\Big{(}e^{(s-t)A_{\mathfrak{c}}}X+z_{\mathfrak{c}}(\theta_{s}\omega)\Big{)}e^{(s-t)A_{\mathfrak{c}}}A_{\mathfrak{c}}X\,\mathrm{d}s,$
@@
$\int_{-\infty}^{t}e^{(t-s)A_{\mathfrak{f}}}\Pi_{\mathfrak{f}}DG_{k}\Big{(}e^{(s-t)A_{\mathfrak{c}}}X+z_{\mathfrak{c}}(\theta_{s}\omega)\Big{)}e^{(s-t)A_{\mathfrak{c}}}A_{\mathfrak{c}}X\,\mathrm{d}s,$
@@
$q(s,\omega)$ $=\int_{t-\tau}^{s}e^{(s-s^{\prime})A_{\mathfrak{f}}}\Pi_{\mathfrak{f}}G_{k}\Big{(}e^{(s^{\prime}-s)A_{\mathfrak{c}}}X+z_{\mathfrak{c}}(s^{\prime},\omega)\Big{)}\,\mathrm{d}s^{\prime}.$
@@
$q_{\tau}(t,X,\omega)=\int_{t-\tau}^{t}e^{(t-s^{\prime})A_{\mathfrak{f}}}\Pi_{\mathfrak{f}}G_{k}\Big{(}e^{(s^{\prime}-t)A_{\mathfrak{c}}}X+z_{\mathfrak{c}}(s^{\prime},\omega)\Big{)}\,\mathrm{d}s^{\prime}.$
@@
$\int_{-\infty}^{t-\tau}e^{(t-s)A_{\mathfrak{f}}}\Pi_{\mathfrak{f}}G_{k}(e^{(s-t)A_{\mathfrak{c}}}X+z_{\mathfrak{c}}(s,\omega))\,\mathrm{d}s.$
The boundedness of $\mathfrak{J}$ implies that the $H_{\mathfrak{f}}$-norm of the integral on the RHS of (115) converges to zero as $\tau$ goes to infinity, and (21) follows.
@@ Appendix B Solving the sACE and its optimal reduced system
For the numerical integration of the stochastic Allen-Cahn equation (76), we adopt a semi-implicit Euler method, where the nonlinearity $-u^{3}$ and the noise term $\,\mathrm{d}{\bm{W}_{t}}$ are treated explicitly, and the linear term $\partial_{x}^{2}u+u$ is treated implicitly; the spatial discretization is handled with a pseudo-spectral method; see [29, Section 6.1] for a detailed description of this scheme applied to a stochastic Burgers-type equation. We have set the time-step size to be $\delta t=10^{-2}$ and the spatial resolution to be $\delta x=L/201$.
@@ C.1 Computation of the double-fold bifurcation diagram
The numerical integration of the stochastic equation (97) is performed in the same way as for the sACE outlined in Appendix B using a semi-implicit Euler method. Here also the nonlinearity and the noise term are treated explicitly, while the linear term is treated implicitly. The spatial discretization is handled with a pseudo-spectral method; see again [29, Section 6.1]. We have set the time-step size to be $\delta t=10^{-2}$ and the spatial resolution to be $\delta x=L/257$. The initial data is set to be $u_{0}=0.5\bm{e}_{1}$ for the results shown in Figs. 12 and 15.
The optimal reduced system (Eq. (LABEL:Eq_reduced_S_shaped)) is integrated with a forward Euler method still with $\delta t=10^{-2}$. The initial condition $J^{n}_{\tau^{*}_{n}}(t_{0};\omega)$ (cf. (100)) for integrating (101) is computed using the trapezoidal rule, and $t_{0}=\max\{\tau_{3}^{*},\tau_{5}^{*}\}$. The initial data for Eq. (LABEL:Eq_reduced_S_shaped) is set to be the projection of the SPDE solution onto $\bm{e}_{1}$ at time $t=t_{0}$.